solve the systems of equations by the substitution method:
5x-4y+=30-5x
3x+6=x+5y+15
I added 5x to both sides to get all variables on one side and got
10x-4y=30 now I not sure where to go from here to solve for y....Please help!
lilchikcadee@msn.com


Answer by unlockmath(1688)

Hello,
Let's set up both equations
5x-4y+=30-5x
3x+6=x+5y+15
like this:
10x-4y=30
-3x-5y=9
Since substitution is required let's set the first equation like this:
x=2/5(y)+3
Now substitute this into the second equation to get:
-3(2/5y+3)-5y=9 Rewritten as:
-6/5(y)-9-5y=9 Add 9 and combine like terms to get:
-19/5 y=18 Multiply both sides by -5/19 to get:
y=-90/31
Plug this into the first to get:
10x-4(-90/31)=30 This works out to:
x=57/31
You can now plug these answers into both and it works out.
Make sense?
RJ
www.math-unlock.com